OpenStudy (anonymous):

you were informed that your auto insurance premium would be increased by 21%(40.95) afer youyr last accident. How much are you paying now after the increase, for insurance/

OpenStudy (anonymous):

235.95

OpenStudy (anonymous):

... beccaaaauuuseee... 40.95 is 21% of 195: (40.95/(21/100)) = (40.95*100)/21 = 195. Aaannddd 195+40.95 = 235.95.